assume that an investor buys 100 shares of stock at 50 per share putting up 70% margin. If the stock rises to $80, how much would the investor have in equity?
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statement showing Computations
Particulars Amount
Price per share                     50.00
Margin = 50*70%                     35.00
Gain per share = 80 - 50                     30.00
Total equity = 30 + 35                     65.00
Investor have in Equity = 65/80 81.25%
Investor have in Equity in $ = 65*100               6,500.00